The following equilibrium constants have been determined for hydrogen sulfide at 25 degree C:
H2S (aq) <=> H^+(aq) + HS^- (aq) ; K'c = 9.510^(-8)
HS^- (aq) <=> H^+ (aq) + S^(2-) (aq) ; K''c = 1.010^(-19)
Calculate the equilibrium constant for the following reaction at the same temperature:
H2S (aq) <=> 2 H^+ (aq) + S^(2-) (aq)
1
Expert's answer
2020-11-14T13:56:15-0500

H2S <=> H+ + HS- K'c = 9.5 ×10-8. ... 1

HS- <=> H+ + S2- K"c = 1.0×10-19 ...2

Add 1 and 2

H2S <=> 2H+ + S2-

K"'c = 9.5×10-8 × 1.0×10-19

K"'c = 9.5 ×10-27